### v1

# Preconditioners for the Stochastic Training of Implicit Neural Representations

Abstract
Implicit neural representations have emerged as a powerful technique for encoding complex continuous multidimensional signals as neural networks, enabling a wide range of applications in computer vision, robotics, and geometry. While Adam is commonly used for training due to its stochastic proficiency, it entails lengthy training durations. To address this, we explore alternative optimization techniques for accelerated training without sacrificing accuracy. Traditional second-order optimizers like L-BFGS are suboptimal in stochastic settings, making them unsuitable for large-scale data sets. Instead, we propose stochastic training using curvature-aware diagonal preconditioners, showcasing their effectiveness across various signal modalities such as images, shape reconstruction, and Neural Radiance Fields (NeRF).
footnotetext: * Equal contribution.footnotetext: Correspondence to: Shin-Fang Chng <shinfang.chng@adelaide.edu.au>, Hemanth Saratchandran <hemanth.saratchandran@adelaide.edu.au>
1 Introduction
Implicit neural representations (INRs) have gained significant traction in the fields of machine learning and computer graphics [57, 60, 12, 28, 10, 56, 35, 55]. Unlike conventional grid-based representations that rely on explicit geometric primitives or parameterized functions [19, 68, 8, 59], INRs take a distinct approach. They define shapes and objects implicitly through trained neural networks. These representations have garnered attention for their ability to efficiently encode complex and highly detailed continuous signals, making them suitable for tasks such as image synthesis [57, 12], shape modeling [35, 56, 11, 38, 1] and robotics [62, 13]. INRs also offer desirable properties, including continuous and differentiable outputs, allowing seamless integration with gradient-based optimizers.
Classically INRs were trained with a ReLU activation, leading to low fidelity reconstructions due to such networks admitting spectral bias [43]. To overcome spectral bias, positional embedding layers (PE) [61] are often added but this increases the parameter count of the network and can lead to noisy gradients. A more recent approach is to use non-traditional activations such as sine [55], Gaussian [44], and wavelets [53]. The use of such activations enable high-frequency encoding without positional embedding layers. The major benefit of these activations over positional embedding layers is their well-behaved gradients [55, 44].
While INRs have opened up new possibilities for realistic and efficient signal representation and manipulation, the training of these architectures still presents a significant research challenge. Curiously, the prevailing approach to training such networks has become almost axiomatic - practitioners typically rely on the Adam optimizer. In the realm of architectures that rely on extensive datasets, the adoption of a stochastic training scheme becomes essential. Surprisingly, the preference for using Adam as the optimizer over a more traditional vanilla optimizer like SGD remains unexplained, despite its widespread usage. In [54], the authors demonstrated remarkable acceleration in training INRs by utilizing second-order optimizers such as L-BFGS [37]. However, an obstacle encountered by [54] was the limited performance of second-order optimizers, such as L-BFGS, in stochastic settings, rendering their methods inefficient for stochastic training of INRs.
In this paper, we investigate the stochastic training of INRs from the perspective of preconditioning. Preconditioners are essential components in optimization and iterative solvers, aiming to enhance convergence and efficiency in numerical methods [50, 37, 5]. They specifically address the condition number, which measures the sensitivity of the solution to small input perturbations. By applying suitable transformations, preconditioners effectively reduce the condition number, facilitating rapid convergence in iterative solvers [16].
We delve into the relationship between the Adam optimizer and preconditioned gradient descent algorithms. Specifically, we demonstrate that Adam can be viewed as employing a diagonal preconditioner based on the Gauss-Newton matrix. This sheds light on why Adam performs remarkably well for INRs compared to vanilla SGD. Furthermore, we propose the use of curvature-aware preconditioners for training INRs, especially for architectures with non-traditional activations such as sine [55], Gaussian [44], or wavelets [53]. We show that employing curvature-aware preconditioners can significantly accelerate training times in stochastic settings.
Our analysis focuses on specific algorithms that leverage these preconditioners, including equilibrated stochastic gradient descent (ESGD) [16], AdaHessian [66], Shampoo [24], and K-FAC [33]. By exploring the potential of these preconditioners, we gain valuable insights into improving the training process for INRs in stochastic settings.
Our main contributions are:
  1. 1.
    An analysis of preconditioning for training INRs unveiling the theoretical underpinnings and illuminating the precise scenarios where preconditioning becomes useful.
  2. 2.
    An empirical validation of our theory by demonstrating the superiority of curvature-aware preconditioning algorithms, such as ESGD, over Adam across diverse signal modalities, including images, shape reconstruction, and neural radiance fields.
2 Related Work
Implicit neural representations (INRs) encode signals as weights within a neural network using low-dimensional coordinates, aiming to preserve smoothness in the outputs [68]. The seminal work of Mildenhall et al. [35] demonstrated the power of INRs by showing their remarkable generalization abilities for novel view synthesis. Since this work, there have been several works on INRs finding applications in image view synthesis, shape reconstruction and robotics [14, 17, 20, 36, 39, 40, 41, 45, 51, 55, 15, 63, 67, 52, 69, 9]. Conventionally these works employ a positional embedding layer to their networks, in order to overcome the spectral bias of a ReLU activation [43]. More recently, several new activations have been employed for INRs, such as sine [55], Gaussian [44], wavelets [53]. These new non-traditional activations are showing superior performance on several vision tasks [55, 15, 44, 53, 65, 32].
Preconditioners were first introduced in the context of numerical linear algebra [50, 22], so as to transform ill-conditioned linear systems into better-conditioned forms, enabling faster convergence of iterative methods. Since then they have found their way into machine learning, where they have been widely employed to enhance the efficiency and effectiveness of optimization algorithms. The use of preconditioning techniques in machine learning can be traced back to early works on optimization, such as the conjugate gradient method and its variants [37]. Over time, researchers have explored various preconditioning strategies tailored to the specific characteristics of machine learning problems, including sparse matrices, non-convex objectives, and large-scale datasets [4, 42, 46, 27]. Several well known optimizers such as ADMM [30], mirror descent [2], quasi-Newton methods [7, 21, 37], adaptive optimizers [16, 66, 18, 24, 33] all employ some form of preconditioning.
3 Analysis
3.1 Preconditioners for Stochastic Training
Preconditioning enhances optimization convergence and efficiency by regulating curvature, particularly in stochastic gradient descent (SGD) [16, 42]. High curvature directions in SGD require a small learning rate to avoid overshooting, leading to slow progress in low curvature directions.
For general objective functions $f$ on parameters $\textbf{x}\in\mathbb{R}^{n}$. Preconditioning is done by employing a linear change of parameters $=D^{\frac{1}{2}}\textbf{x}$ x, where in practise $D^{\frac{1}{2}}$ is a non-singular matrix. With this change of parameters a new objective function $\widetilde{f}$, defined on $\widetilde{\textbf{x}}$, can be defined as
$\widetilde{f}(\widetilde{\textbf{x}})=f(\textbf{x}).$ (1)
Denoting the gradient of the original objective function by $\nabla f$ and its Hessian by $H(f)=\nabla^{2}f(\textbf{x})$. The chain rule gives
$\nabla\widetilde{f}(\widetilde{\textbf{x}})$ $=D^{-\frac{1}{2}}\nabla f$ (2)
$H(\widetilde{f})(\widetilde{\textbf{x}})$ $=(D^{-\frac{1}{2}})^{T}(H(f)(\textbf{x}))(D^{-\frac{1}{2}}).$ (3)
Equation (2) exhibits the key idea of training with a preconditioner. If we consider gradient descent on the transformed parameter space, a gradient descent iteration takes the form
$\widetilde{\textbf{x}}_{t}=\widetilde{\textbf{x}}_{t-1}-\eta\nabla_{\widetilde{\textbf{x}}_{t-1}}\widetilde{f}$ (4)
Applying equations (1) and (2) this corresponds to doing the update
$\textbf{x}_{t}=\textbf{x}_{t-1}-\eta D^{-\frac{1}{2}}\nabla_{\textbf{x}_{t-1}}f.$ (5)
The main endeavour of preconditioning for optimization is to find a preconditioning matrix $D$ such that the new Hessian, given by equation (3), has equal curvature in all directions. This will allow the update (5) to move faster along small curvature directions, leading to faster convergence. Algorithm 1 gives the basic pseudocode for preconditioned SGD.
Algorithm 1 Preconditioned SGD
initial point $x_{0}$, iterations $N$, learning rate $\eta$, preconditioner $D$.
for $t=0,\cdots,N$ do
     $g_{t}\leftarrow\nabla_{x_{t}}f$
     $D_{t}\leftarrow D(x_{t})$
     $x_{t+1}\leftarrow x_{t}-\eta D(x_{t})^{-\frac{1}{2}}g_{t}$
end for
From equation (3) we see that the best preconditioner to use would be the absolute value of the Hessian itself. In the case of a positive definite Hessian, this is precisely what Newton’s method does [37]. The main issue with this preconditioner is that the cost of storing and inverting the preconditioner will be $\mathcal{O}(n^{3})$ for an objective function of $n$ parameters. This will be extremely costly for the training of overparameterized neural networks.
We list some of the main preconditioners that are being used in the literature.
Diagonal Preconditioners:
are matrices that represent some form of a diagonal approximation of the Hessian. Examples include the Jacobi preconditioner, which is given by $D^{J}=|diag(H)|$, where ||$\cdot$|| ⋅ | is element-wise absolute value and $H$ is the Hessian. In practical settings, the absolute value is taken so that the preconditioner is positive semidefinite. The equilibration preconditioner is a diagonal matrix which has the 2-norm of each row (or column) on the diagonal: $D^{E}=||H_{i,\cdot}||$. Diagonal preconditioners are often used because of their ability to be stored and inverted in linear time. Diagonal preconditioners are also known as adapative learning rates in the literature.
The algorithm Equilibrated Stochastic Gradient Descent (ESGD) [16] employs $D^{E}$ as a preconditioner to condition stochastic gradient descent. In [16], it is shown that applying ESGD leads to faster training times for autoencoder tasks when compared to vanilla SGD, RMSProp [48] and a preconditioned SGD using $D^{J}$ as preconditioner. The pseudocode for ESGD is given in algorithm 2.
Algorithm 2 ESGD
initial point $x_{0}$, iterations $N$, learning rate $\eta$.
$D\leftarrow 0$
for $t=0,\cdots,N$ do
     $g_{t}\leftarrow\nabla_{x_{t}}f$
     $D_{t}\leftarrow D_{t-1}+D^{E}_{t}$
     $x_{t+1}\leftarrow x_{t}-\eta D_{t-1}^{-\frac{1}{2}}g_{t}$
end for
To efficiently calculate $D^{E}$ during each iteration, it is unnecessary to compute the entire Hessian and then determine the 2-norm of each row, which would consume $\mathcal{O}(n^{2})$ memory for an objective function with $n$ parameters. Alternatively, one leverages the following matrix free estimator result of Bradley and Murray [6]
$||H_{i,\cdot}||=\mathbb{E}[(Hv)^{2}]\text{ for }v\sim\mathcal{N}(0,1).$ (6)
Eqn (6) allows one to compute $D^{J}$ at each iteration with a cost equivalent to one backpropagation.
AdaHessian [66] is a recent algorithm that applies a diagonal preconditioner in the setting of moving averages and thus can be seen as a way of enhancing the Adam optimizer [26]. Given an objective function $f$, AdaHessian, like Adam and RMSProp, computes a first order moment $m_{t}$ and a second order moment $v_{t}$:
$m_{t}$ $=\frac{(1-\beta_{1})\sum_{i=1}^{t}\beta_{1}^{t-i}\textbf{$g_{i}$}}{1-\beta_{1}^{t}}$ (7)
$v_{t}$ $=\bigg{(}\sqrt{\frac{(1-\beta_{2})\sum_{i=1}^{t}D_{i}D_{i}}{1-\beta_{2}^{t}}}\bigg{)}^{k}$ (8)
where $g_{i}$ denotes the gradient iteration $i$, $D_{i}$ denotes the diagonal of the Hessian at iteration $i$ and $0<\beta_{1}$, $\beta_{2}<1$ are hyperparameters that are also used in the Adam optimizer. When $k=1$, equation (8) defines a moving average preconditioner that uses the diagonal of the Hessian, which implies in such cases AdaHessian is using the Jacobi preconditioner $D^{J}$. Algorithm 3 gives the pseudocode for AdaHessian.
Algorithm 3 AdaHessian
initial point $x_{0}$, iterations $N$, learning rate $\eta$, exponential decay rates $\beta_{1}$, $\beta_{2}$, power $k$
$v_{0}\leftarrow 0$
$m_{0}\leftarrow 0$
for $t=1,\cdots,N$ do
     $g_{t}\leftarrow\nabla_{x_{t}}f$
     $D_{t}\leftarrow diag(H_{t})=\text{diagonal of current Hessian}$
     Update $v_{t}$ and $m_{t}$ using eqns (7), (8)
     $x_{t+1}\leftarrow x_{t}-\eta\frac{m_{t}}{v_{t}}$
end for
To efficiently calculate $D^{J}$ during each iteration, one can use the following matrix free estimator result of Bekas et al. [3]
$Diag(H)=\mathbb{E}[v\odot Hv]\text{ for }v\sim\mathcal{R}$ (9)
where $\mathcal{R}$ represents a Rademacher distribution.
Kronecker Factored Preconditioners:
are preconditioners of the form $D=A^{T}A$, where $A=A_{1}\otimes\cdots\otimes A_{k}$ and $\otimes$ denotes the Kronecker product. The algorithm K-FAC [33] employs a Kronecker factored preconditioner.
Others: There are several second order optimisation algorithms that use preconditioners that do not fall into the above three categories such as, The Gauss-Newton method, The Conjugate Gradient method, Limited memory Broyden-Fisher-Golfarb- Shanno (L-BFGS) [37]. It was shown in [54] that these algorithms do not train well in stochastic settings. Another algorithm of importance for this work is Shampoo, which applies a left and right preconditioning matrix, see [24] for details.
3.2 Condition Number
In Sec. 3.1, we explained that SGD will move faster through the loss landscape when the curvatures in each direction are similar. Quantitatively the regularity of the loss landscape at a point is measured via the condition number of the Hessian.
Definition 1.
Given a matrix $A$ we define the condition number of A by
$\kappa(A):=\frac{\sigma_{\max}(A)}{\sigma_{\min}(A)}.$ (10)
Using eqn (3) we seek a preconditioner $D$ such that $\kappa((D^{-\frac{1}{2}})^{T}(H(f)(\textbf{x}))(D^{-\frac{1}{2}}))$ is closer to 1 than $\kappa(H)$. This implies that the preconditioner $D$ is regularising the curvatures of the landscape to be equal in each direction.
In [23], Guggenheimer et al. prove the following upper bound on the conditioner number of an $N\times N$ Hessian matrix.
$\kappa(H)<\frac{2}{|det(H)|}\bigg{(}\frac{||H||_{F}}{\sqrt{N}}\bigg{)}.$ (11)
Furthermore, they show that a tight bound is expected when all the singular values, except for the smallest, are roughly equal. In [16], it is shown that by applying an equilibrated preconditioner the above bound can be tightened by a factor of $det(D^{E})\bigg{(}\frac{||H||_{F}}{\sqrt{N}}\bigg{)}^{N}$.
In our empirical validation (based on [16]), we examined a one-dimensional INR scenario. The INR consisted of two hidden layers, each with 16 neurons, aimed at fitting the signal $f(x)=\sin(2\pi x)+\sin(6\pi x)$. During optimization using gradient descent, we employed an equilibrated preconditioner $D^{E}$ and a Jacobi preconditioner $D^{J}$. At each optimization step, we computed both the original Hessian of the gradient descent algorithm and the transformed Hessian using equation (3). Fig. 1 illustrates the experimental results, revealing that the equilibrated descent algorithm exhibits a trajectory with a significantly lower-conditioned Hessian compared to the Jacobi preconditioner and the original Hessian of a gradient descent.
(a) Figure 1: Comparison of condition number (10) before and after applying equilbrated and Jacobi preconditioners on a 1D signal regression. The condition number of the equilbrated preconditioner has significantly reduced (solid blue vs. dotted blue) than Jacobi (solid red vs. dotted red)
(b) Figure 2: Comparison of training convergence (PSNR) between Adam and ESGD on a NeRF task with ReLU-PE. The ReLU-PE INR trained with Adam significantly outperformed the one trained with ESGD, suggesting curvature aware preconditioners offer no benefit for ReLU-PE.
Figure 3: Visualization of the sparsity patterns in the Hessian-vector product matrices for the penultimate layer of ReLU-PE (Left) and Gaussian networks (Right): white pixels represent zero values, and black pixels denotes non-zero values.
3.3 Why INRs train so well with Adam compared to SGD
Given an objective function $L(\theta)=|f(\theta)|^{2}$ for a non-linear function $f$. The Jacobian $J_{L}$ of $L$ can be computed via the chain rule: $J_{L}=2J_{f}^{T}f$, where $J_{f}$ denotes the Jacobian of $f$. The Hessian of $L$, $H_{L}$, can be computed by one more application of the chain rule
$H_{L}=2J_{f}^{T}J_{f}+2H_{f}f.$ (12)
We can then approximate the Hessian $H_{L}$ by the term $2J_{f}^{T}J_{f}$. This leads to
$H_{L}\approx 2J_{f}^{T}J_{f}=\frac{1}{2|f|^{2}}J_{L}^{T}J_{L}=\frac{1}{2L}J_{L}^{T}J_{L}.$ (13)
Eqn. 13 yields the basic idea of the Gauss-Newton method. By taking the Gauss-Newton matrix $J_{L}^{T}J_{L}$ as an approximation to the Hessian $H_{L}$, we can perform an optimization update that uses the step: $\Delta_{GN}=-\frac{1}{J^{T}J}\nabla L$. The Gauss-Newton matrix is in general only positive semi-definite. Therefore, when implementing the algorithm one often adds a small damping factor $\lambda$ so the implemented step is $\Delta_{GN}=-\frac{1}{J^{T}J+\lambda}\nabla L$. A similar derivation can be made for more general loss functions used in non-linear least squares type regression problems [37]. It follows that the Gauss-Newton method can be seen as a preconditioned SGD that uses a Hessian approximate preconditioner given by the Gauss-Newton matrix $J^{T}J$.
The Adam optimizer involves two key steps. The first is to compute a first moment estimate of the form $m_{t}=\beta_{1}m_{t-1}+(1-\beta_{1})g_{t}$, where $g_{t}$ is the gradient of a given objective function $f$ at iteration $t$, and $m_{0}=0$. The second is to compute a second moment estimate $v_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})g_{t}^{2}$, where $g_{t}^{2}=g_{t}\odot g_{t}$, $\odot$ denotes the Hadamard product, and $v_{0}=0$. The numbers $10<\beta_{1}<\beta_{2}<$ 1 are hyperparameters. The algorithm then applies a bias correction to the above first and second moment estimates, and then takes a moving average to perform an update step based on $-\frac{m_{t}}{\sqrt{v_{t}}}$, see [26] for details.
For the following discussion, we will only focus on analyzing the second raw moment estimate, as this is related to the Gauss-Newton matrix. Given an objective function $f$, observe that we can express $g_{t}^{2}=g_{t}\odot g_{t}=Diag(J^{T}J)$, where $J$ denotes the Jacobian of $f$. Thus the second raw moment estimate of Adam can be recast as:
$v_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})Diag(J_{t}^{T}J_{t})$ (14)
where $J_{t}$ denotes the Jacobian at iteration $t$. Thus we see that Adam applies a preconditioner given by the diagonal of the Gauss-Newton matrix. Since the Gauss-Newton matrix can be seen as a first order approximation to the Hessian, Adam can be seen as using curvature information to adapt the learning rate for faster optimization. This suggests that Adam can outperform SGD which employs no preconditioner.
We empirically verified this insight in Fig. 4 on a binary occupancy and NeRF experiment. The details of the architectures and experiments are given in Sec. 4.2 and 4.3. It is clear from this figure that Adam outperformed SGD in both signal modalities. Similar results hold in the case of a ReLU/ReLU-PE INR, see supp. material.
3.4 Activation matters for curvature preconditioning
In Sec. 3.1, we highlighted that many implemented preconditioners in practice do not directly compute a Hessian matrix. Instead, they rely on approximations through Hessian vector products. Consequently, comprehending the Hessian vector product’s structure becomes crucial for INRs.
Theorem 3.1.
Let $F$ denote a sine- or Gaussian-activated INR. Let $\mathcal{L}(\theta)$ denote the MSE loss associated to $F$ and a training set $(X,Y)$. Let $H$ denote the Hessian of $\mathcal{L}$ at a fixed parameter point $\theta$. Then given a non-zero vector $v$ the Hessian vector product $Hv$ will be a dense vector.
While it may seem rather costly for a sine- or Gaussian-activated INR to have dense Hessian vector products when applied to the MSE loss function, we point out that the cost is much less than storing a dense Hessian.
Theorem 3.2.
Let $F$ denote a ReLU/ReLU-PE-activated INR. Let $\mathcal{L}(\theta)$ denote the MSE loss associated to $F$ and a training set $(X,Y)$. Let $H$ denote the Hessian of $\mathcal{L}$ at a fixed parameter point $\theta$. Then given a non-zero vector $v$ the Hessian vector product $Hv$ will be a sparse vector.
Theorems 3.1 and 3.2 also hold for the binary cross-entropy loss.
Theorem 3.2 implies that when applying a preconditioner, built from a Hessian vector product, to a ReLU/ReLU-PE INR, such as ESGD [16] or AdaHessian [66], only a few components of the gradient will be scaled during the update step. This means that the preconditioner will not be adding much to the optimization process. On the other hand, theorem 3.1 implies that applying a curvature preconditioner to the gradient of the MSE loss of a sine/Gaussian-activated INR should help optimization as many of the gradient components will get scaled by a quantity depending on the curvature of the loss landscape.
Fig. 3 shows the results of a Hessian vector product of the MSE loss of a ReLU-PE and Gaussian INR with respect to the weights in the penultimate layer, see Sec. 4.1 for the network configuration. The random vector was chosen from a Rademacher distribution. Hessian vector product with respect to weights in the other layers of the neural network exhibited similar properties. As can be seen from the experiment the ReLU-PE network produces a sparse vector, as implied by theorem 3.2, and the Gaussian one produces a dense vector, as implied by theorem 3.1.
Theorem 3.1 suggests that training a ReLU/ReLU-PE INR with a curvature aware preconditioner that uses Hessian vector products would not lead to any significant advantage. In order to empirically verify this, we trained a ReLU-PE INR on a NeRF task, see Sec. 4.3 for details on the configuration, using ESGD and Adam. As shown in Fig. 2, the ReLU-PE INR trained with Adam significantly outperforms the one trained with ESGD. This suggests that for ReLU/ReLU-PE activated INRs, such curvature aware preconditioners offer no benefit.
Figure 4: Comparison of training convergence between Adam and SGD on optimizing a 3D occupancy field (Left) and a NeRF (Right). Binary cross-entropy (BCE) loss measures the accuracy of the shape reconstruction, while PSNR loss assesses the scene fidelity.
Figure 5: Comparison of training convergence in Mean Squared Error (MSE) loss (a) and computational complexity of various preconditioner (b and c). ESGD demonstrates superior performance compared to other curvature-aware preconditioners on an image reconstruction task, striking a balance between accuracy and computational efficiency.
4 Experiments
We compare different preconditioners in training of INR on various popular benchmarks [25, 55, 54, 31]: 2D image reconstruction, 3D shape reconstruction and novel view synthesis using NeRF. We focus on analyzing Gaussian networks, with variance denoted by $\sigma$, with preconditioners, which are ESGD [16] [^1] https://github.com/crowsonkb/esgd, AdaHessian [66] [^2] https://github.com/amirgholami/adahessian, Shampoo [24] [^3] https://github.com/google-research/google-research/tree/master/scalable_shampoo/pytorch, and Kronecker-based preconditioner [33] [^4] https://github.com/Thrandis/EKFAC-pytorch, see supp. material for analysis for other non-traditional activations and the reproducibility details. Remark: We observed that in the implementation of AdaHessian, an equilbrated preconditioner $D^{E}$ is utilized, even though the paper primarily focuses on using a Jacobi preconditioner [16]. Therefore, we introduced a Jacobi preconditioner $D^{J}$ into AdaHessian, which we refer to as AdaHessian(J). We use the term AdaHessian(E) for the original implementation. Since the local Hessian computed with minibatches is generally noisy [66], we opted to recompute the $D^{E}$ for every $N$ iterations for computational efficiency, we use $N=50$ for all the experiments. Additionally, we also employed a moving average technique [66, 26] to mitigate the impact of noise.
4.1 Image Reconstruction
Given pixel coordinates $\mathbf{x}\in\mathbb{R}^{2}$, we optimized an INR to regress the associated RGB values $c\in\mathbb{R}^{3}$. We use a 5-layer Gaussian-activated network with 256 neurons and $\sigma=0.05$. The network is trained stochastically using mini-batches of size $512$. As seen in Fig. 5(a), while both AdaHessian(J) and AdaHessian(E) demonstrate faster convergence than ESGD in terms of iterations (Fig. 5b), it is important to note that their training time per iteration is longer than ESGD.
4.2 3D Shape Reconstruction
Given 3D point clouds $\mathbf{x}\in\mathbb{R}^{3}$, we optimize an INR to represent the occupancy field of 3D shapes.  [^5] http://graphics.stanford.edu/data/3Dscanrep/ For training, we sampled 12 million 3D points – one-third of the points were sampled uniformly within the volume, and remaining two-thirds of the points were sampled near the mesh surface and perturbed with random Gaussian noise using sigma of 0.1 and 0.01, respectively [25]. We used a 5-layer Gaussian network with 256 neurons and $\sigma=0.09$. The network was trained stochastically using mini-batches of size $50000$. We optimized a binary cross-entropy loss between the predicted occupancy and the true occupancy. Fig. 6(a) shows that ESGD has superior convergence than other preconditioner-type optimizers, see more results in supp. material. Interestingly, when dealing with higher mode signals, AdaHessian(J) and AdaHessian(E) no longer showed superior convergence, as in Sec. 4.1. We speculated that the local Hessians might become noisy in higher mode signals, thereby impacted the gradient updates. However, this noise did not have a significant impact on ESGD’s convergence since the preconditioner updates were performed every $N$ iterations.
Figure 6: Comparison of training convergence (a) in average of Binary Cross- Entropy (BCE) loss of various preconditioners on optimizing a 3D occupancy field. We show the intermediate results (b) at epoch 500 for Adam [26](Left) and ESGD [16](Right). Compared to INR trained with Adam, ESGD has reconstructed the shapes with significantly improved fidelities (zoomed in for better visibility).
4.3 NeRF
NeRF is a compelling strategy for utilizing an INR to model 3D objects and scenes using multi-view 2D images. This approach shows promise for generating high-fidelity reconstruction in novel view synthesis task [35, 47, 15, 29]. Given 3D points $\mathbf{x}\in\mathbb{R}^{3}$ and viewing direction, NeRF estimates the radiance field of a 3D scene which maps each input 3D coordinate to its corresponding volume density $\sigma\in\mathbb{R}$ and directional emitted color $\mathbf{c}\in\mathbb{R}^{3}$ [35, 29, 15]. We used a 8-layer Gaussian-activated network with 256 neurons and $\sigma=0.1$. Table 1 shows that ESGD achieves on average faster convergence than Adam, see more results in the supp. material.
Adam ESGD
Scene Train Iteration Time Test Iteration Time Test
PSNR $\downarrow$ ($s$$\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ $\downarrow$ ($s$$\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
fern 25.38 200K 368.43 24.38 0.74 0.28 120K 270.81 24.41 0.74 0.30
flower 29.17 200K 469.31 25.67 0.78 0.12 120K 263.93 25.65 0.78 0.13
orchids 23.3 200K 501.90 19.56 0.61 0.24 120K 269.01 19.74 0.61 0.21
fortress 30.18 140K 333.70 28.93 0.81 0.15 140K 334.56 28.6 0.80 0.17
leaves 21.36 200K 482.34 19.56 0.61 0.24 120K 271.25 19.43 0.59 0.25
Table 1: Quantitative results of NeRF on instances from the LLFF dataset [34]. Note that the report time solely refers to the optimization update step.
5 Conclusion
We investigated the use of preconditioners in stochastic training of INRs. We found that employing a curvature-aware preconditioner significantly speeds up training for INRs admitting non-traditional activations such as Gaussian functions. Experimental validation across signal modalities, including images, shape reconstruction, and neural radiance fields, using ESGD and AdaHessian algorithms supported our findings. However, for INRs with ReLU/ReLU-PE activations, curvature-aware preconditioners offered no significant advantage over Adam. This was confirmed in a comparison study on a NeRF task, where Adam outperformed ESGD for a ReLU-PE INR.
6 Limitations
Our investigation into condition numbers, as detailed in Sec. 3.2, revealed that an equilibrated preconditioner effectively reduces the condition number of the loss landscape compared to a Jacobi preconditioner. This theoretical insight was established in [16] and further validated through empirical experiments involving autoencoder tasks. Surprisingly, our image regression experiment (see Figure 5) demonstrated that a Jacobi preconditioner yielded faster training results than an equilibrated preconditioner. These findings suggest that when employing a preconditioned gradient descent algorithm for optimizing implicit neural representations, the influence of the condition number on convergence may be more complex than initially anticipated. Further research in this area is warranted to gain a deeper understanding.

### v2

@@
-Preconditioners for the Stochastic Training of Implicit Neural Representations
+Preconditioners for the Stochastic Training of Neural Fields
 Abstract
-Implicit neural representations have emerged as a powerful technique for encoding complex continuous multidimensional signals as neural networks, enabling a wide range of applications in computer vision, robotics, and geometry.
-While Adam is commonly used for training due to its stochastic proficiency, it entails lengthy training durations.
-To address this, we explore alternative optimization techniques for accelerated training without sacrificing accuracy.
-Traditional second-order optimizers like L-BFGS are suboptimal in stochastic settings, making them unsuitable for large-scale data sets.
-Instead, we propose stochastic training using curvature-aware diagonal preconditioners, showcasing their effectiveness across various signal modalities such as images, shape reconstruction, and Neural Radiance Fields (NeRF).
-footnotetext: * Equal contribution.footnotetext: Correspondence to: Shin-Fang Chng <shinfang.chng@adelaide.edu.au>, Hemanth Saratchandran <hemanth.saratchandran@adelaide.edu.au>
+Neural fields encode continuous multidimensional signals as neural networks, enabling diverse applications in computer vision, robotics, and geometry.
+While Adam is effective for stochastic optimization, it often requires long training times.
+To address this, we explore alternative optimization techniques to accelerate training without sacrificing accuracy.
+Traditional second-order methods like L-BFGS are unsuitable for stochastic settings.
+We propose a theoretical framework for training neural fields with curvature-aware diagonal preconditioners, demonstrating their effectiveness across tasks such as image reconstruction, shape modeling, and Neural Radiance Fields (NeRF)  [^1] Source code is available at https://github.com/sfchng/preconditioner_neural_fields.git..
+Figure 1: We demonstrate that curvature-aware preconditioners can significantly accelerate training convergence for neural fields with suitable activations.
+Specifically, ESGD (a curvature-aware preconditioned gradient descent algorithm [14]) improves convergence for wavelet, sine, and Gaussian activations, while Adam performs better for ReLU network with positional encoding (ReLU(PE)).
+In a 2D image reconstruction task (Fig.
+1(a)), a wavelet-based neural field trained with ESGD achieves 29.36 PSNR (Fig.
+1(e)) at the $10^{\text{th}}$ epoch, outperforming Adam’s 26.61 PSNR (Fig.
+1(d)).
+In a 3D binary occupancy task (Fig.
+1(b)), ESGD enables faster convergence for sine activations, as shown by the improved armadillo mesh reconstruction at the $1000^{\text{th}}$ epoch (Fig.
+1(g) vs.
+Fig.
+1(f)).
+In a Neural Radiance Fields (NeRF) task (Fig.
+1(c)), a Gaussian-based model trained with ESGD achieves lower reconstruction error (Fig.
+1(i), $0.0695$) compared to Adam (Fig.
+1(h), $0.0743$), as seen in the error map (darker areas indicate lower error; zoom in 4×4$\times$4 × for clarity).
+11footnotetext: Equal contribution. Correspondence to: Shin-Fang Chng $<$shinfangchng@gmail.com$>$, Hemanth Saratchandran $<$hemanth.saratchandran.adelaide.edu.au$>$.
 1 Introduction
-Implicit neural representations (INRs) have gained significant traction in the fields of machine learning and computer graphics [57, 60, 12, 28, 10, 56, 35, 55].
-Unlike conventional grid-based representations that rely on explicit geometric primitives or parameterized functions [19, 68, 8, 59], INRs take a distinct approach.
-They define shapes and objects implicitly through trained neural networks.
-These representations have garnered attention for their ability to efficiently encode complex and highly detailed continuous signals, making them suitable for tasks such as image synthesis [57, 12], shape modeling [35, 56, 11, 38, 1] and robotics [62, 13].
-INRs also offer desirable properties, including continuous and differentiable outputs, allowing seamless integration with gradient-based optimizers.
-Classically INRs were trained with a ReLU activation, leading to low fidelity reconstructions due to such networks admitting spectral bias [43].
-To overcome spectral bias, positional embedding layers (PE) [61] are often added but this increases the parameter count of the network and can lead to noisy gradients.
-A more recent approach is to use non-traditional activations such as sine [55], Gaussian [44], and wavelets [53].
-The use of such activations enable high-frequency encoding without positional embedding layers.
-The major benefit of these activations over positional embedding layers is their well-behaved gradients [55, 44].
-While INRs have opened up new possibilities for realistic and efficient signal representation and manipulation, the training of these architectures still presents a significant research challenge.
-Curiously, the prevailing approach to training such networks has become almost axiomatic - practitioners typically rely on the Adam optimizer.
-In the realm of architectures that rely on extensive datasets, the adoption of a stochastic training scheme becomes essential.
-Surprisingly, the preference for using Adam as the optimizer over a more traditional vanilla optimizer like SGD remains unexplained, despite its widespread usage.
-In [54], the authors demonstrated remarkable acceleration in training INRs by utilizing second-order optimizers such as L-BFGS [37].
-However, an obstacle encountered by [54] was the limited performance of second-order optimizers, such as L-BFGS, in stochastic settings, rendering their methods inefficient for stochastic training of INRs.
-In this paper, we investigate the stochastic training of INRs from the perspective of preconditioning.
-Preconditioners are essential components in optimization and iterative solvers, aiming to enhance convergence and efficiency in numerical methods [50, 37, 5].
-They specifically address the condition number, which measures the sensitivity of the solution to small input perturbations.
-By applying suitable transformations, preconditioners effectively reduce the condition number, facilitating rapid convergence in iterative solvers [16].
-We delve into the relationship between the Adam optimizer and preconditioned gradient descent algorithms.
-Specifically, we demonstrate that Adam can be viewed as employing a diagonal preconditioner based on the Gauss-Newton matrix.
-This sheds light on why Adam performs remarkably well for INRs compared to vanilla SGD.
-Furthermore, we propose the use of curvature-aware preconditioners for training INRs, especially for architectures with non-traditional activations such as sine [55], Gaussian [44], or wavelets [53].
-We show that employing curvature-aware preconditioners can significantly accelerate training times in stochastic settings.
-Our analysis focuses on specific algorithms that leverage these preconditioners, including equilibrated stochastic gradient descent (ESGD) [16], AdaHessian [66], Shampoo [24], and K-FAC [33].
-By exploring the potential of these preconditioners, we gain valuable insights into improving the training process for INRs in stochastic settings.
+In recent years, neural fields, also known as implicit neural representations, have garnered significant attention in machine learning and computer graphics [55, 57, 9, 25, 8, 53, 32, 54].
+Unlike traditional grid-based representations [17, 64, 6, 56], neural fields implicitly encode shapes and objects through neural networks.
+These representations are highly efficient for capturing complex, high-resolution signals, making them well-suited for tasks such as image synthesis [55, 9], shape modeling [32, 53], and robotics [59, 10].
+Classically, neural fields trained with ReLU activations suffer from spectral bias and limited fidelity [40].
+To address this, positional encoding (PE) layers [58] were introduced but at the cost of increased model complexity and noisy gradients.
+More recent work has explored non-standard activations, such as sine [54], Gaussian [41] and wavelets [49] which enable high-frequency encoding without positional encodings and produce well-behaved gradients [50].
+While neural fields have enabled realistic and efficient representations of continuous signals, progress on minimizing the training time of such representations remains a significant research challenge.
+Currently, the Adam optimizer is the de facto algorithm for training neural fields, particularly in applications involving datasets where a stochastic training regime is essential.
+Yet, the preference for Adam over simpler alternatives such as stochastic gradient descent (SGD) has not been rigorously justified for neural fields, despite its widespread adoption.
+Notably, [50] demonstrated substantial gains in training speed for neural fields using second-order optimizers such as L-BFGS [34].
+However, second-order methods are often impractical for stochastic settings [50], limiting their effectiveness for large-scale data, where stochastic training is essential.
+In this paper, we approach neural field training from the perspective of preconditioning.
+Preconditioners improve convergence by reducing the condition number of the optimization problem [46, 34, 4].
+We develop a theoretical framework linking Adam’s performance to its implicit use of a diagonal preconditioner derived from the Gauss-Newton matrix and proceed to identify conditions under which curvature-aware diagonal preconditioners, such as those in ESGD (equilibrated stochastic gradient descent) [14], enhance training for neural fields with sine, Gaussian, or wavelet activations.
+For ReLU activations, we show that such preconditioning provides limited benefit.
+Our experiments validate the proposed framework across a range of tasks, including 2D image reconstruction, 3D binary occupancy, and Neural Radiance Fields (NeRF).
+As demonstrated in Fig.
+1, ESGD consistently outperforms Adam for neural fields with wavelet, sine, and Gaussian activations, which, as predicted by our theory, is due to their better-conditioned Hessians.
+In contrast, Adam performs better for ReLU-PE activations, which are associated with poorly conditioned Hessians.
+Furthermore, we evaluated other curvature-aware algorithms, including ESGD, AdaHessian [62], Shampoo [21], and K-FAC [30], offering valuable insights into their effectiveness in stochastic training settings.
 Our main contributions are:
   1. 1.
-    An analysis of preconditioning for training INRs unveiling the theoretical underpinnings and illuminating the precise scenarios where preconditioning becomes useful.
+    We present a theoretical framework for preconditioning in neural field training, uncovering key conditions and highlighting when preconditioning is beneficial.
   2. 2.
-    An empirical validation of our theory by demonstrating the superiority of curvature-aware preconditioning algorithms, such as ESGD, over Adam across diverse signal modalities, including images, shape reconstruction, and neural radiance fields.
+    We empirically validate our framework, demonstrating the superiority of curvature-aware algorithms like ESGD over Adam across diverse tasks, including image reconstruction, 3D shape modeling, and NeRF.
 2 Related Work
-Implicit neural representations (INRs) encode signals as weights within a neural network using low-dimensional coordinates, aiming to preserve smoothness in the outputs [68].
-The seminal work of Mildenhall et al.
-[35] demonstrated the power of INRs by showing their remarkable generalization abilities for novel view synthesis.
-Since this work, there have been several works on INRs finding applications in image view synthesis, shape reconstruction and robotics [14, 17, 20, 36, 39, 40, 41, 45, 51, 55, 15, 63, 67, 52, 69, 9].
-Conventionally these works employ a positional embedding layer to their networks, in order to overcome the spectral bias of a ReLU activation [43].
-More recently, several new activations have been employed for INRs, such as sine [55], Gaussian [44], wavelets [53].
-These new non-traditional activations are showing superior performance on several vision tasks [55, 15, 44, 53, 65, 32].
-Preconditioners were first introduced in the context of numerical linear algebra [50, 22], so as to transform ill-conditioned linear systems into better-conditioned forms, enabling faster convergence of iterative methods.
-Since then they have found their way into machine learning, where they have been widely employed to enhance the efficiency and effectiveness of optimization algorithms.
-The use of preconditioning techniques in machine learning can be traced back to early works on optimization, such as the conjugate gradient method and its variants [37].
-Over time, researchers have explored various preconditioning strategies tailored to the specific characteristics of machine learning problems, including sparse matrices, non-convex objectives, and large-scale datasets [4, 42, 46, 27].
-Several well known optimizers such as ADMM [30], mirror descent [2], quasi-Newton methods [7, 21, 37], adaptive optimizers [16, 66, 18, 24, 33] all employ some form of preconditioning.
-3 Analysis
-3.1 Preconditioners for Stochastic Training
+Neural Fields
+represent signals as weights within a neural network, using low-dimensional coordinates to maintain smoothness in the output [64].
+The influential work by Mildenhall et al.
+[32] showcased the potential of neural fields for novel view synthesis, demonstrating their impressive generalization capabilities.
+Since then, neural fields have seen widespread application across image view synthesis, shape reconstruction, and robotics [11, 15, 18, 33, 35, 36, 38, 42, 47, 54, 12, 60, 63, 48, 65, 7].
+Traditionally, these approaches integrate a positional embedding layer to counteract the spectral bias introduced by ReLU activations [40].
+However, recent work has explored alternative, non-standard activations such as sine [54], Gaussian [41], wavelets [49], and sinc [51].
+These activations have shown to outperform conventional approaches in various vision tasks, delivering enhanced performance in capturing high-frequency details and reducing reliance on positional embeddings [54, 12, 41, 49, 61, 29].
+Preconditioners
+were initially introduced in numerical linear algebra [46, 20] to transform ill-conditioned linear systems into well-conditioned forms, thereby accelerating the convergence of iterative solvers.
+Over time, preconditioning techniques have become integral to machine learning, where they enhance the efficiency and performance of optimization algorithms.
+Early applications in machine learning can be traced to foundational optimization methods, such as the conjugate gradient method and its variants [34].
+As machine learning problems have grown more complex—often involving sparse matrices, non-convex objectives, and large-scale datasets—researchers have developed specialized preconditioning strategies to meet these unique challenges [3, 39, 43, 24].
+Many widely-used optimizers incorporate preconditioning, including methods like ADMM [27], mirror descent [2], quasi-Newton methods [5, 19, 34], and adaptive optimizers [14, 62, 16, 21, 30].
+These preconditioned optimizers have become essential for addressing the complex landscapes of modern machine learning models, driving improvements in both convergence speed and solution quality.
+3 Notation
+The neural fields we consider in this paper will be defined by a depth $L$ neural network with layer widths $\{n_{1},\ldots,n_{L}\}$.
+We let $X\in\mathbb{R}^{N\times n_{0}}$ denote the training data, with $n_{0}$ being the dimension of the input and $N$ being the number of training samples.
+The output at layer $k$ will be denoted by $F_{k}$ and is defined by
+$F_{k}=\begin{cases}F_{L-1}W_{L}+b_{L},&k=L\\ \phi(F_{k-1}W_{k}+b_{k}),&k\in[L-1]\\ X,&k=0\end{cases}$ (1)
+where the weights $W_{k}\in\mathbb{R}^{n_{k-1}\times n_{k}}$ and the biases $b_{k}\in\mathbb{R}^{n_{k}}$ and $\phi$ is an activation applied component wise.
+The notation $[m]$ is defined by $[m]=\{1,\ldots,m\}$.
+We will denote the loss function that we train the neural fields with by $\mathcal{L}$.
+In many experiments in the paper this will be the MSE loss [37].
+Our theoretical results will be primarily for the case where $\phi$ is given by one of the activation sine, Gaussian, wavelet or ReLU.
+For more details on how these activations are used within the context of neural fields we refer the reader to [54, 41, 49].
+4 Theoretical Framework
+4.1 Overview of Preconditioners
 Preconditioning enhances optimization convergence and efficiency by regulating curvature, particularly in stochastic gradient descent (SGD) [14, 39].
@@
 end for
-From equation (3) we see that the best preconditioner to use would be the absolute value of the Hessian itself.
+From Eq.
+4 we see that the best preconditioner to use would be the absolute value of the Hessian itself.
 In the case of a positive definite Hessian, this is precisely what Newton’s method does [34].
-The main issue with this preconditioner is that the cost of storing and inverting the preconditioner will be $\mathcal{O}(n^{3})$ for an objective function of $n$ parameters.
-This will be extremely costly for the training of overparameterized neural networks.
+The main issue with this preconditioner is that the cost of storing and inverting the preconditioner will be $\mathcal{O}(n^{3})$ for an objective function of $n$ parameters [34].
+This will be extremely costly for the training of overparameterized neural networks whose Hessians are extremely large.
 We list some of the main preconditioners that are being used in the literature.
@@
 The equilibration preconditioner is a diagonal matrix which has the 2-norm of each row (or column) on the diagonal: $D^{E}=||H_{i,\cdot}||$.
+The preconditioners $D^{J}$ and $D^{E}$ are known as curvature-aware preconditioners as they seek to give a good approximation to the Hessian.
 Diagonal preconditioners are often used because of their ability to be stored and inverted in linear time.
@@
 In [14], it is shown that applying ESGD leads to faster training times for autoencoder tasks when compared to vanilla SGD, RMSProp [45] and a preconditioned SGD using $D^{J}$ as preconditioner.
-The pseudocode for ESGD is given in algorithm 2.
-Algorithm 2 ESGD
-initial point $x_{0}$, iterations $N$, learning rate $\eta$.
-$D\leftarrow 0$
-for $t=0,\cdots,N$ do
-     $g_{t}\leftarrow\nabla_{x_{t}}f$
-     $D_{t}\leftarrow D_{t-1}+D^{E}_{t}$
-     $x_{t+1}\leftarrow x_{t}-\eta D_{t-1}^{-\frac{1}{2}}g_{t}$
-end for
-To efficiently calculate $D^{E}$ during each iteration, it is unnecessary to compute the entire Hessian and then determine the 2-norm of each row, which would consume $\mathcal{O}(n^{2})$ memory for an objective function with $n$ parameters.
-Alternatively, one leverages the following matrix free estimator result of Bradley and Murray [6]
-$||H_{i,\cdot}||=\mathbb{E}[(Hv)^{2}]\text{ for }v\sim\mathcal{N}(0,1).$ (6)
-Eqn (6) allows one to compute $D^{J}$ at each iteration with a cost equivalent to one backpropagation.
 AdaHessian [62] is a recent algorithm that applies a diagonal preconditioner in the setting of moving averages and thus can be seen as a way of enhancing the Adam optimizer [23].
-Given an objective function $f$, AdaHessian, like Adam and RMSProp, computes a first order moment $m_{t}$ and a second order moment $v_{t}$:
-$m_{t}$ $=\frac{(1-\beta_{1})\sum_{i=1}^{t}\beta_{1}^{t-i}\textbf{$g_{i}$}}{1-\beta_{1}^{t}}$ (7)
-$v_{t}$ $=\bigg{(}\sqrt{\frac{(1-\beta_{2})\sum_{i=1}^{t}D_{i}D_{i}}{1-\beta_{2}^{t}}}\bigg{)}^{k}$ (8)
-where $g_{i}$ denotes the gradient iteration $i$, $D_{i}$ denotes the diagonal of the Hessian at iteration $i$ and $0<\beta_{1}$, $\beta_{2}<1$ are hyperparameters that are also used in the Adam optimizer.
-When $k=1$, equation (8) defines a moving average preconditioner that uses the diagonal of the Hessian, which implies in such cases AdaHessian is using the Jacobi preconditioner $D^{J}$.
-Algorithm 3 gives the pseudocode for AdaHessian.
-Algorithm 3 AdaHessian
-initial point $x_{0}$, iterations $N$, learning rate $\eta$, exponential decay rates $\beta_{1}$, $\beta_{2}$, power $k$
-$v_{0}\leftarrow 0$
-$m_{0}\leftarrow 0$
-for $t=1,\cdots,N$ do
-     $g_{t}\leftarrow\nabla_{x_{t}}f$
-     $D_{t}\leftarrow diag(H_{t})=\text{diagonal of current Hessian}$
-     Update $v_{t}$ and $m_{t}$ using eqns (7), (8)
-     $x_{t+1}\leftarrow x_{t}-\eta\frac{m_{t}}{v_{t}}$
-end for
-To efficiently calculate $D^{J}$ during each iteration, one can use the following matrix free estimator result of Bekas et al. [3]
-$Diag(H)=\mathbb{E}[v\odot Hv]\text{ for }v\sim\mathcal{R}$ (9)
-where $\mathcal{R}$ represents a Rademacher distribution.
 Kronecker Factored Preconditioners:
@@
 It was shown in [50] that these algorithms do not train well in stochastic settings.
-Another algorithm of importance for this work is Shampoo, which applies a left and right preconditioning matrix, see [24] for details.
 4.2 Condition Number
 In Sec.
-3.1, we explained that SGD will move faster through the loss landscape when the curvatures in each direction are similar.
-Quantitatively the regularity of the loss landscape at a point is measured via the condition number of the Hessian.
-Definition 1.
-Given a matrix $A$ we define the condition number of A by
-$\kappa(A):=\frac{\sigma_{\max}(A)}{\sigma_{\min}(A)}.$ (10)
-Using eqn (3) we seek a preconditioner $D$ such that $\kappa((D^{-\frac{1}{2}})^{T}(H(f)(\textbf{x}))(D^{-\frac{1}{2}}))$ is closer to 1 than $\kappa(H)$.
-This implies that the preconditioner $D$ is regularising the curvatures of the landscape to be equal in each direction.
-In [23], Guggenheimer et al. prove the following upper bound on the conditioner number of an $N\times N$ Hessian matrix.
-$\kappa(H)<\frac{2}{|det(H)|}\bigg{(}\frac{||H||_{F}}{\sqrt{N}}\bigg{)}.$ (11)
-Furthermore, they show that a tight bound is expected when all the singular values, except for the smallest, are roughly equal.
-In [16], it is shown that by applying an equilibrated preconditioner the above bound can be tightened by a factor of $det(D^{E})\bigg{(}\frac{||H||_{F}}{\sqrt{N}}\bigg{)}^{N}$.
-In our empirical validation (based on [16]), we examined a one-dimensional INR scenario.
-The INR consisted of two hidden layers, each with 16 neurons, aimed at fitting the signal $f(x)=\sin(2\pi x)+\sin(6\pi x)$.
-During optimization using gradient descent, we employed an equilibrated preconditioner $D^{E}$ and a Jacobi preconditioner $D^{J}$.
-At each optimization step, we computed both the original Hessian of the gradient descent algorithm and the transformed Hessian using equation (3).
-Fig.
-1 illustrates the experimental results, revealing that the equilibrated descent algorithm exhibits a trajectory with a significantly lower-conditioned Hessian compared to the Jacobi preconditioner and the original Hessian of a gradient descent.
-(a) Figure 1: Comparison of condition number (10) before and after applying equilbrated and Jacobi preconditioners on a 1D signal regression.
-The condition number of the equilbrated preconditioner has significantly reduced (solid blue vs.
-dotted blue) than Jacobi (solid red vs.
-dotted red)
-(b) Figure 2: Comparison of training convergence (PSNR) between Adam and ESGD on a NeRF task with ReLU-PE.
-The ReLU-PE INR trained with Adam significantly outperformed the one trained with ESGD, suggesting curvature aware preconditioners offer no benefit for ReLU-PE.
-Figure 3: Visualization of the sparsity patterns in the Hessian-vector product matrices for the penultimate layer of ReLU-PE (Left) and Gaussian networks (Right): white pixels represent zero values, and black pixels denotes non-zero values.
-3.3 Why INRs train so well with Adam compared to SGD
+4.1, we explained that SGD will move faster through the loss landscape when the curvatures in each direction are similar.
+The regularity of the loss landscape at a point is quantified via the condition number of the Hessian.
+Definition 4.1.
+Given a matrix $A$ we define the condition number $\kappa$ of A by
+$\kappa(A):=\frac{\sigma_{\max}(A)}{\sigma_{\min}(A)},$ (7)
+where $\sigma$ is the eigenvalue.
+Using Eq.
+4 we observe that if we take a preconditioner $D$ such that $\kappa((D^{-\frac{1}{2}})^{T}(H(f)(\textbf{x}))$ $(D^{-\frac{1}{2}}))$ is closer to 1 than $\kappa(H)$, then this implies that the preconditioner $D$ is regularising the curvatures of the landscape to be equal in each direction, which yields a better training landscape for a gradient descent based algorithm [34].
+To empirically illustrate the concept of preconditioning, we examined a one-dimensional curve fitting scenario.
+Specifically, we considered a feedforward network with two hidden layers, each containing 16 neurons, tasked with fitting the signal $f(x)=\sin(2\pi x)+\sin(6\pi x)$ using a Gaussian activation [41].
+During optimization with gradient descent, we applied two types of preconditioners: an equilibrated preconditioner $D^{E}$ and a Jacobi preconditioner $D^{J}$.
+At each optimization step, we computed both the original Hessian associated with gradient descent and the transformed Hessian using Eq.
+4.
+As shown in Fig.
+2, the results demonstrate that applying either the Jacobi or equilibrated preconditioner consistently reduced the condition number of the Hessian compared to the unconditioned case, highlighting the effectiveness of preconditioning in improving optimization.
+Figure 2: Fig.
+2(a) compares the condition number (7) on a 1D signal regression task before (dotted line) and after (solid line) applying equilbrated $D^{E}$ and Jacobi $D^{J}$ preconditioners.
+The equilibrated preconditioner significantly reduces the Hessian’s condition number, as shown by the larger gap between the yellow solid and dotted lines compared to the green solid and dotted lines representing the Jacobi preconditioner, thus achieves faster convergence as seen in Fig.
+2(b).
+4.3 Training with Adam
+Figure 3: Comparison of training convergence for neural fields with different activations – ReLU with positional encoding (ReLU (PE)), wavelet, sine and Gaussian – using Adam and SGD optimizers across three tasks: 2D image reconstruction (Fig.
+3(a)), 3D binary occupancy reconstruction (Fig.
+3(b)), and NeRF (Fig.
+3(c)).
+In all cases, the Adam optimizer outperforms SGD, indicating that leveraging curvature information facilitates faster convergence.
+In this section, we aim to explain why training with Adam is generally more efficient than training with SGD for a variety of neural fields.
+While this observation is well-known within the community, we include it here to provide context and motivation for the theoretical framework developed in Sec.
+4.4.
 Given an objective function $L(\theta)=|f(\theta)|^{2}$ for a non-linear function $f$.
 The Jacobian $J_{L}$ of $L$ can be computed via the chain rule: $J_{L}=2J_{f}^{T}f$, where $J_{f}$ denotes the Jacobian of $f$.
-The Hessian of $L$, $H_{L}$, can be computed by one more application of the chain rule
+The Hessian of $L$, $H_{L}$, can be computed by an application of the chain rule
 $H_{L}=2J_{f}^{T}J_{f}+2H_{f}f.$ (8)
-We can then approximate the Hessian $H_{L}$ by the term $2J_{f}^{T}J_{f}$. This leads to
+We can then approximate the Hessian $H_{L}$ by the term $2J_{f}^{T}J_{f}$. Leading to the first order approximation
 $H_{L}\approx 2J_{f}^{T}J_{f}=\frac{1}{2|f|^{2}}J_{L}^{T}J_{L}=\frac{1}{2L}J_{L}^{T}J_{L}.$ (9)
-Eqn.
+Eq.
 9 yields the basic idea of the Gauss-Newton method.
@@
 A similar derivation can be made for more general loss functions used in non-linear least squares type regression problems [34].
-It follows that the Gauss-Newton method can be seen as a preconditioned SGD that uses a Hessian approximate preconditioner given by the Gauss-Newton matrix $J^{T}J$.
+It follows that the Gauss-Newton method can be seen as a preconditioned SGD that uses a first order (first derivatives) Hessian approximate preconditioner given by the Gauss-Newton matrix $J^{T}J$.
+Using the above observations we now show how to view Adam as an optimizer that applies a diagonal preconditioner.
 The Adam optimizer involves two key steps.
@@
 Thus we see that Adam applies a preconditioner given by the diagonal of the Gauss-Newton matrix.
-Since the Gauss-Newton matrix can be seen as a first order approximation to the Hessian, Adam can be seen as using curvature information to adapt the learning rate for faster optimization.
+Since the Gauss-Newton matrix can be seen as a first order approximation to the Hessian, Adam can be seen as using an approximation to curvature to adapt the learning rate for faster optimization.
 This suggests that Adam can outperform SGD which employs no preconditioner.
-We empirically verified this insight in Fig.
-4 on a binary occupancy and NeRF experiment.
-The details of the architectures and experiments are given in Sec.
-4.2 and 4.3.
-It is clear from this figure that Adam outperformed SGD in both signal modalities.
-Similar results hold in the case of a ReLU/ReLU-PE INR, see supp.
-material.
-3.4 Activation matters for curvature preconditioning
+We empirically validated this insight across three distinct neural field tasks: image reconstruction, binary occupancy, and NeRF.
+For each task, we compared the performance of ReLU(PE) activated networks trained with both SGD and Adam against networks using wavelet, sine, and Gaussian activations.
+As shown in Fig.
+3, Adam consistently outperforms SGD in all cases.
+This suggests that Adam’s use of a preconditioner significantly enhances its performance for neural field applications, see Sec.
+5 for the details of the architectures and experimental setup.
+4.4 Main theorems
+Figure 4: Percentage sparsity of Hessian-vector product matrices across all layers throughout the training process of a 2D image reconstruction task, using ESGD to train networks with different activations.
+The x-axis ranges from 00 to $1$, with values closer to 1 indicating higher sparsity.
 In Sec.
-3.1, we highlighted that many implemented preconditioners in practice do not directly compute a Hessian matrix.
-Instead, they rely on approximations through Hessian vector products.
-Consequently, comprehending the Hessian vector product’s structure becomes crucial for INRs.
+4.3, we discussed how Adam can be viewed as applying a diagonal preconditioner based on the diagonal of the Gauss-Newton matrix, which uses first derivatives to approximate the Hessian matrix.
+Here, we aim to identify the architectural properties that enable effective training of neural fields with diagonal preconditioners that give a better approximation to the Hessian.
+Our primary theorems demonstrate that the choice of activation function is crucial for enabling efficient training with such preconditioners, highlighting its central role in achieving effective curvature-aware optimization for neural fields.
+Recall Sec.
+4.1 that diagonal preconditoners such as the equilibrated precondition $D^{E}$ or the Jacobi preconditioner $D^{J}$ all involve the Hessian matrix.
+In the context of neural networks the Hessian of the loss landscape is often an extremely large matrix.
+Thus in practical settings one often replaces the Hessian with a Hessian vector product which is much cheaper to store [34].
+Our first theorem shows that the Hessian vector product of the loss landscape associated to a sine, Gaussian, wavelet or sinc activated neural field produces a dense vector.
 Theorem 4.2.
-Let $F$ denote a sine- or Gaussian-activated INR.
+Let $F$ denote a sine, Gaussian, wavelet or sinc activated neural field.
 Let $\mathcal{L}(\theta)$ denote the MSE loss associated to $F$ and a training set $(X,Y)$.
@@
 Theorem 4.3.
-Let $F$ denote a ReLU/ReLU-PE-activated INR.
+Let $F$ denote a ReLU/ReLU(PE)-activated neural field.
 Let $\mathcal{L}(\theta)$ denote the MSE loss associated to $F$ and a training set $(X,Y)$.
@@
 Then given a non-zero vector $v$ the Hessian vector product $Hv$ will be a sparse vector.
-Theorems 3.1 and 3.2 also hold for the binary cross-entropy loss.
-Theorem 3.2 implies that when applying a preconditioner, built from a Hessian vector product, to a ReLU/ReLU-PE INR, such as ESGD [16] or AdaHessian [66], only a few components of the gradient will be scaled during the update step.
+Theorem 4.2 and Theorem 4.3 also hold for the binary cross-entropy loss. The proofs of Theorem 4.2 and Theorem 4.3 are given in the supp.
+Theorem 4.3 implies that when applying a preconditioner, built from a Hessian vector product, to a ReLU/ReLU(PE) INR, such as ESGD [14] or AdaHessian [62], only a few components of the gradient will be scaled during the update step.
 This means that the preconditioner will not be adding much to the optimization process.
-On the other hand, theorem 3.1 implies that applying a curvature preconditioner to the gradient of the MSE loss of a sine/Gaussian-activated INR should help optimization as many of the gradient components will get scaled by a quantity depending on the curvature of the loss landscape.
+On the other hand, Theorem 4.2 implies that applying a preconditioner built from the Hessian to the gradient of the MSE loss of a sine, Gaussian and wavelet-activated neural field should help optimization as many of the gradient components will get scaled by a quantity depending on the curvature of the loss landscape.
+This yields a clear difference between ReLU/ReLU(PE) activated neural fields and those activated with a sine, Gaussian, wavelet activation.
+In particular, Theorem 4.2 implies that neural fields admitting a sine, Gaussian, wavelet activation have the ability to be trained efficiently with a diagonal curvature aware preconditioner such as the equilibrated preconditioner $D^{E}$ defined in Sec.
+4.1.
 Fig.
-3 shows the results of a Hessian vector product of the MSE loss of a ReLU-PE and Gaussian INR with respect to the weights in the penultimate layer, see Sec.
-4.1 for the network configuration.
-The random vector was chosen from a Rademacher distribution.
-Hessian vector product with respect to weights in the other layers of the neural network exhibited similar properties.
-As can be seen from the experiment the ReLU-PE network produces a sparse vector, as implied by theorem 3.2, and the Gaussian one produces a dense vector, as implied by theorem 3.1.
-Theorem 3.1 suggests that training a ReLU/ReLU-PE INR with a curvature aware preconditioner that uses Hessian vector products would not lead to any significant advantage.
-In order to empirically verify this, we trained a ReLU-PE INR on a NeRF task, see Sec.
-4.3 for details on the configuration, using ESGD and Adam.
-As shown in Fig.
-2, the ReLU-PE INR trained with Adam significantly outperforms the one trained with ESGD.
-This suggests that for ReLU/ReLU-PE activated INRs, such curvature aware preconditioners offer no benefit.
-Figure 4: Comparison of training convergence between Adam and SGD on optimizing a 3D occupancy field (Left) and a NeRF (Right).
-Binary cross-entropy (BCE) loss measures the accuracy of the shape reconstruction, while PSNR loss assesses the scene fidelity.
-Figure 5: Comparison of training convergence in Mean Squared Error (MSE) loss (a) and computational complexity of various preconditioner (b and c).
-ESGD demonstrates superior performance compared to other curvature-aware preconditioners on an image reconstruction task, striking a balance between accuracy and computational efficiency.
+4 shows the results of a Hessian vector product of the MSE loss of a ReLU(PE) and Gaussian neural field for an image reconstruction task for the weights in all layers of the network layer throughout training.
+The random vector was chosen from a Rademacher distribution and the reported figures are given as the mean of the Hessian of all layers averaged over all iterations of training.
+As demonstrated by the experiment, the ReLU(PE) network generates a sparse Hessian-vector product, as predicted by Theorem 4.3, while the Gaussian network produces a dense vector, as indicated by Theorem 4.2.
+Furthermore, as we will observe in Sec.
+5, this sparsity helps explain why Adam performs well for neural fields with ReLU (ReLU(PE)) activations, whereas an algorithm like ESGD [14] does not.
+This is because Adam does not rely on Hessian-vector product.
 5 Experiments
-We compare different preconditioners in training of INR on various popular benchmarks [25, 55, 54, 31]: 2D image reconstruction, 3D shape reconstruction and novel view synthesis using NeRF.
-We focus on analyzing Gaussian networks, with variance denoted by $\sigma$, with preconditioners, which are ESGD [16] [^1] https://github.com/crowsonkb/esgd, AdaHessian [66] [^2] https://github.com/amirgholami/adahessian, Shampoo [24] [^3] https://github.com/google-research/google-research/tree/master/scalable_shampoo/pytorch, and Kronecker-based preconditioner [33] [^4] https://github.com/Thrandis/EKFAC-pytorch, see supp.
-material for analysis for other non-traditional activations and the reproducibility details.
-Remark: We observed that in the implementation of AdaHessian, an equilbrated preconditioner $D^{E}$ is utilized, even though the paper primarily focuses on using a Jacobi preconditioner [16].
+Figure 5: Comparison of training convergence and computational complexity for various preconditioners.
+We evaluated a Gaussian-activated neural field on the lion instance from the DIV2K dataset.
+ESGD demonstrates superior convergence compared to other preconditioners, striking a balance between accuracy and computational efficiency.
+Note: We provide the time-based comparisons in the supp.
+(Sec 3.1).
+Similar analysis for other activations and additional instances from the DIV2K dataset are available in the supp.
+(Sec.
+3.2).
+We compared various preconditioners over popular benchmarks [22, 54, 50, 28, 52]: 2D image reconstruction, 3D shape reconstruction and novel view synthesis with NeRF.
+Our primary focus was on curvature-aware preconditoners that are well-suited for the stochastic training of neural network.
+Specifically, we compared the following methods: ESGD [14] [^2] https://github.com/crowsonkb/esgd, AdaHessian [62] [^3] https://github.com/amirgholami/adahessian, Shampoo [21] [^4] https://github.com/google-research/google-research/tree/master/scalable_shampoo/pytorch, and Kronecker-based preconditioner [30] [^5] https://github.com/Thrandis/EKFAC-pytorch.
+We provide the details of each algorithm including the reproducibility details of each experiment and machines we used to run each experiment in the supp.
+In the main paper, we focused on analyzing Gaussian networks, see supp.
+for additional analysis for other non-traditional activations and other neural field tasks.
+Remark for AdaHessian
+We observed that in the implementation of AdaHessian, an equilbrated preconditioner $D^{E}$ is utilized, even though the paper primarily focuses on using a Jacobi preconditioner [14].
 Therefore, we introduced a Jacobi preconditioner $D^{J}$ into AdaHessian, which we refer to as AdaHessian(J).
 We use the term AdaHessian(E) for the original implementation.
-Since the local Hessian computed with minibatches is generally noisy [66], we opted to recompute the $D^{E}$ for every $N$ iterations for computational efficiency, we use $N=50$ for all the experiments.
-Additionally, we also employed a moving average technique [66, 26] to mitigate the impact of noise.
+Remark for ESGD
+We further modified the ESGD by applying the squared Hessian diagonal preconditioner in two settings: using an exponential moving average (ESGD) and using an inifinity norm (ESGD-max) [23].
+We also opted to recompute the $D^{E}$ for every $N$ iterations for computational efficiency, we use $N=100$ for all the experiments.
 5.1 Image Reconstruction
-Given pixel coordinates $\mathbf{x}\in\mathbb{R}^{2}$, we optimized an INR to regress the associated RGB values $c\in\mathbb{R}^{3}$.
-We use a 5-layer Gaussian-activated network with 256 neurons and $\sigma=0.05$.
-The network is trained stochastically using mini-batches of size $512$.
-As seen in Fig.
-5(a), while both AdaHessian(J) and AdaHessian(E) demonstrate faster convergence than ESGD in terms of iterations (Fig.
-5b), it is important to note that their training time per iteration is longer than ESGD.
-4.2 3D Shape Reconstruction
-Given 3D point clouds $\mathbf{x}\in\mathbb{R}^{3}$, we optimize an INR to represent the occupancy field of 3D shapes.
-[^5] http://graphics.stanford.edu/data/3Dscanrep/ For training, we sampled 12 million 3D points – one-third of the points were sampled uniformly within the volume, and remaining two-thirds of the points were sampled near the mesh surface and perturbed with random Gaussian noise using sigma of 0.1 and 0.01, respectively [25].
-We used a 5-layer Gaussian network with 256 neurons and $\sigma=0.09$.
-The network was trained stochastically using mini-batches of size $50000$.
-We optimized a binary cross-entropy loss between the predicted occupancy and the true occupancy.
-Fig.
-6(a) shows that ESGD has superior convergence than other preconditioner-type optimizers, see more results in supp.
-material.
-Interestingly, when dealing with higher mode signals, AdaHessian(J) and AdaHessian(E) no longer showed superior convergence, as in Sec.
-4.1.
-We speculated that the local Hessians might become noisy in higher mode signals, thereby impacted the gradient updates.
-However, this noise did not have a significant impact on ESGD’s convergence since the preconditioner updates were performed every $N$ iterations.
-Figure 6: Comparison of training convergence (a) in average of Binary Cross- Entropy (BCE) loss of various preconditioners on optimizing a 3D occupancy field.
-We show the intermediate results (b) at epoch 500 for Adam [26](Left) and ESGD [16](Right).
-Compared to INR trained with Adam, ESGD has reconstructed the shapes with significantly improved fidelities (zoomed in for better visibility).
-4.3 NeRF
-NeRF is a compelling strategy for utilizing an INR to model 3D objects and scenes using multi-view 2D images.
-This approach shows promise for generating high-fidelity reconstruction in novel view synthesis task [35, 47, 15, 29].
-Given 3D points $\mathbf{x}\in\mathbb{R}^{3}$ and viewing direction, NeRF estimates the radiance field of a 3D scene which maps each input 3D coordinate to its corresponding volume density $\sigma\in\mathbb{R}$ and directional emitted color $\mathbf{c}\in\mathbb{R}^{3}$ [35, 29, 15].
-We used a 8-layer Gaussian-activated network with 256 neurons and $\sigma=0.1$.
-Table 1 shows that ESGD achieves on average faster convergence than Adam, see more results in the supp.
-material.
+We evaluated on a 2D image reconstruction task, where given pixel coordinates $\mathbf{x}\in\mathbb{R}^{2}$, we optimized a neural network to regress the corresponding RGB values $c\in\mathbb{R}^{3}$.
+We used a 5-layer Gaussian-activated network with 256 neurons.
+We used $\sigma=0.05$ for the Gaussian activation.
+We trained the network using mini-batches of size $512$.
+We compared the training convergence of different preconditioners and the Adam optimizer on the lion instance from the DIV2K dataset [1] in Fig.
+5(a).
+Additional results on the rest of the DIV2K dataset are provided in the supp.
+As predicted by Theorem 4.2, Gaussian network equipped with the dense Hessian-vector product exhibits faster convergence when optimized with curvature-aware preconditioned methods compared to Adam.
+Notably, ESGD demonstrates faster convergence and has comparable training time and memory usage per iteration compared to Adam.
+ESGD and AdaHessian obtain Hessian information by computing Hessian-vector products, which makes them very computational efficient.
+Each Hessian-vector product is generally as fast as a single gradient computation, as demonstrated in Fig.
+5(b).
+In contrast, other preconditioners, particularly Shampoo are more computationally expensive because they involve Kronecker product computations for preconditioner matrices, see supp.
+for details.
 Adam ESGD
@@
 flower 29.17 200K 469.31 25.67 0.78 0.12 120K 263.93 25.65 0.78 0.13
+horns 24.95 60K 130.82 21.82 0.70 0.34 120K 227.58 20.78 0.63 0.44
 orchids 23.3 200K 501.90 19.56 0.61 0.24 120K 269.01 19.74 0.61 0.21
+room 31.20 110K 194.05 31.60 0.93 0.12 140K 201.88 30.52 0.90 0.20
 fortress 30.18 140K 333.70 28.93 0.81 0.15 140K 334.56 28.6 0.80 0.17
 leaves 21.36 200K 482.34 19.56 0.61 0.24 120K 271.25 19.43 0.59 0.25
-Table 1: Quantitative results of NeRF on instances from the LLFF dataset [34]. Note that the report time solely refers to the optimization update step.
+trex 25.50 80K 164.96 22.04 0.71 0.31 120K 209.25 22.21 0.75 0.23
+Table 1: Quantitative results of NeRF on all instances from the LLFF dataset [31].
+The reported time solely refers to the optimization update step.
+Note: Results for the Blender synthetic [32] and BLEFF [60] datasets are available in supp.
+5.2 3D Binary Occupancy Reconstruction
+We evaluated on learning a binary occupancy field, where given 3D point coordinates $\mathbf{x}\in\mathbb{R}^{3}$, we optimized a neural network to represent the occupancy field of 3D shapes.
+Following existing works [22], we sampled one million 3D points for training – one-third of the points were sampled uniformly within the volume, and the remaining two-thirds of the points were sampled near the mesh surface and perturbed with random Gaussian noise using sigma of 0.1 and 0.01, respectively [22].
+We trained the network using binary cross-entropy (BCE) loss, which compares the predicted occupancy and the true occupancy.
+We used a 5-layer Gaussian network with 256 neurons and $\sigma=0.09$.
+We trained the network using mini-batches of size $50000$.
+We compared the training convergence of different preconditioners and the Adam optimizer on the armadillo instance from the Stanford 3D Scanning dataset.
+[^6] http://graphics.stanford.edu/data/3Dscanrep/ Fig.
+6 shows that ESGD exhibits superior convergence than other preconditioner-type optimizers.
+We showcase a qualitative result comparing the mesh reconstruction of ESGD and Adam given the same number of iteration in Fig.
+7.
+Interestingly, when dealing with higher mode signals, both AdaHessian and Shampoo no longer demonstrated superior convergence, unlike in the 2D image reconstruction task in Fig.
+5(a).
+We speculate that the local Hessians may become noisy with higher mode signals, thereby impacting the gradient updates.
+In contrast, this noise did not significantly affect ESGD’s convergence because its preconditioner updates were performed every $N$ iterations.
+Hence, we will focus on comparing Adam and ESGD for the experiments in the following section.
+Figure 6: Comparison of training convergence for various preconditioners.
+We evaluated a Gaussian-activated neural field on a 3D binary occupancy reconstruction task on the armadillo instance.
+ESGD demonstrates superior convergence compared to other preconditioners, striking a balance between accuracy and computational efficiency.
+Note: Similar analysis for other activations are available in the supp.
+Figure 7: Qualitative results for the dragon instance from Stanford 3D Scanning dataset.
+Compared to the Gaussian-based neural field trained with Adam (Fig.
+7(a)), ESGD (Fig.
+7(b)) has reconstructed the shapes with significantly improved fidelity at epoch 500 (zoom in 4×4$\times$4 × for better visibility).
+Note: Qualitative results for the rest of the instances from Stanford 3D Scanning Dataset are available in supp.
+5.3 Neural Radiance Fields (NeRF)
+Neural Radiance Fields (NeRF) have recently emerged as a pioneering approach in novel view synthesis, utilizing neural networks to model volumetric fields of 3D scenes using multi-view 2D images [32, 44, 12, 26, 13].
+Specifically, given 3D points $\mathbf{x}\in\mathbb{R}^{3}$ and viewing direction, NeRF estimates the 5D radiance field of a 3D scene which maps each input 3D coordinate to its corresponding volume density $\sigma\in\mathbb{R}$ and directional emitted color $\mathbf{c}\in\mathbb{R}^{3}$ [32, 26, 12].
+We used an 8-layer Gaussian-activated network with 256 neurons and $\sigma=0.1$.
+Recall that in the previous section we demonstrated that ESGD exhibited superior performance in terms of convergence and computational complexity compared to other preconditioners.
+Therefore, we focused on comparing Adam and ESGD here.
+We determined the same training PSNR at which both model converged and reported the corresponding optimization metrics, including the number of iterations and the total time taken for the update computation.
+We evaluated the PSNR on a hold-out test scene.
+Table 1.
+shows that ESGD achieves on average faster convergence than Adam, see more results in the supp.
+6 Limitations
+While our work theoretically established the suitability of curvature-aware preconditioners for training neural fields with non-traditional activations such as sine, Gaussian, and wavelets, it also highlighted their ineffectiveness for neural fields employing ReLU-PE activations.
+Specifically, we showed that Adam performs well for ReLU-PE due to its use of a diagonal preconditioner derived from the Gauss-Newton matrix, rather than relying on Hessian vector products.
+However, this raises an important limitation of our work: we did not explore whether alternative preconditioners, potentially more effective than the Gauss-Newton matrix, could be developed to improve the training of ReLU-PE activated neural fields beyond what Adam achieves.
+This is particularly relevant since ReLU-PE remains widely used in practical applications.
+This question represents an avenue for future research, which we leave it as future work.
 7 Conclusion
-We investigated the use of preconditioners in stochastic training of INRs.
-We found that employing a curvature-aware preconditioner significantly speeds up training for INRs admitting non-traditional activations such as Gaussian functions.
-Experimental validation across signal modalities, including images, shape reconstruction, and neural radiance fields, using ESGD and AdaHessian algorithms supported our findings.
-However, for INRs with ReLU/ReLU-PE activations, curvature-aware preconditioners offered no significant advantage over Adam.
-This was confirmed in a comparison study on a NeRF task, where Adam outperformed ESGD for a ReLU-PE INR.
-6 Limitations
-Our investigation into condition numbers, as detailed in Sec.
-3.2, revealed that an equilibrated preconditioner effectively reduces the condition number of the loss landscape compared to a Jacobi preconditioner.
-This theoretical insight was established in [16] and further validated through empirical experiments involving autoencoder tasks.
-Surprisingly, our image regression experiment (see Figure 5) demonstrated that a Jacobi preconditioner yielded faster training results than an equilibrated preconditioner.
-These findings suggest that when employing a preconditioned gradient descent algorithm for optimizing implicit neural representations, the influence of the condition number on convergence may be more complex than initially anticipated.
-Further research in this area is warranted to gain a deeper understanding.
+We explored the impact of preconditioners on the stochastic training of neural fields, developing a theoretical framework based on Hessian-vector products.
+Our analysis demonstrates that curvature-aware preconditioners can significantly accelerate training for neural fields with non-traditional activations, such as sine, Gaussian, and wavelet.
+Experimental validation across various tasks—including image reconstruction, shape modeling, and neural radiance fields—using ESGD and AdaHessian confirmed these benefits.
+In contrast, for neural fields with ReLU or ReLU-PE activations, our theory predicted no significant advantage from curvature-aware preconditioning, highlighting the activation-dependent nature of these methods.
+For practitioners employing activations like sine, Gaussian, or wavelet, robust curvature-aware algorithms provide an efficient approach for stochastic training across diverse applications.